OpenStudy (anonymous):

What is the equation of a line passing through (3, 4) and having a slope of -5? A) y = 5x + 19 B) y = -5x + 11 C) y = -5x + 19 D) y = -1/5x + 19 E) y = -1/5x + 11 @nikato @jdoe0001 @Mertsj

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (nikato):

Do you know the point slope form? If you do, plug in the given info in it and solve for y

OpenStudy (anonymous):

No I don't, what's the form? Lol.

OpenStudy (nikato):

Oh it's that^. And m=slope (x1,y1) are your points

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 3}}\quad ,&{\color{blue}{ 4}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -5 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green} m}(x-{\color{red}{ x_1}})\qquad \textit{plug in your values and solve for "y"}\)

OpenStudy (anonymous):

y-4=-5*x-3 ? Is that right?

OpenStudy (jdoe0001):

\(\bf y-{\color{blue}{ y_1}}={\color{green} m}(x-{\color{red}{ x_1}}) \) <---- point-slope form

OpenStudy (anonymous):

the form is y-y1=m(x-x1) m=slope=-5 y1=4 x1=3

OpenStudy (jdoe0001):

y-4=-5* (x-3)

OpenStudy (jdoe0001):

don't forget to distribute the -5

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

y=-5y + 19, right?

OpenStudy (jdoe0001):

y=-5y + 19, right? \(\large \checkmark\)

OpenStudy (anonymous):

Thanks :)